Working on a DE Tales from the Crypt that is having an odd switch matrix issue. 2 switches in column 7 (switch 52 - Super VUK and switch 55 - Power Scoop) do not register a majority of the time as tested directly at the board with a jumper wire directly at CN8 & CN10 with those connectors from the playfield removed. All the other switches register every time you activate them with the jumper wire but 52 and 55 only register occasionally.
I've removed/reflowed the solder for the header pins at both CN8 and CN10. I tested Q49 and it was good but I replaced it anyways. I've tested the bussed resistor networks at RA6 and RA11 and they are good. I also swapped out the 74LS244 at 8J and the 6821 at 8H even tho both of those tested fine with my chip tester. RA16 tests good as well.
In looking at the schematics, switch 52 at CN10 pin 6 goes to the 4011 at 5J and switch 55 at CN10 pin 2 goes to the 4011 at 6J before the signal goes to the PIA, which I've already replaced. Is it possible that both 4011's have an isolated issue causing these two switches to not register most of the time?